OpenStudy (anonymous):

what are the solutions to the equation 2x 2 +9x=5 (btw its supposed to be 2x squared)

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

2+11x=5

OpenStudy (anonymous):

x = 5/13

OpenStudy (anonymous):

11-5 then 2/-6 then its -3 for x

OpenStudy (anonymous):

is this correct

OpenStudy (luffingsails):

Hi maddy, best to write in proper form first, \[2x^{2} +9x = 5\] \[2x^{2}+9x-5 = 0\] Now attempt to factor: \[(2x-1)(x+5) = 0\] Then set each term equal to zero, and you will have your solutions! 2x-1 = 0; x = 1/2 x+5 = 0; x = -5 Hope that helps!

OpenStudy (anonymous):

i used the x method

OpenStudy (anonymous):

Maddy, use the ^ sign to show something is squared Sorrye my answer is wrong because I did now the other 2 was squared. so next time write like 2x^2 + 9x =5 use the ^

OpenStudy (luffingsails):

@142343, I think you missed the squared on the x... that changes the method entirely. Hope it is clearer now.